OpenStudy (anonymous):

Find the work done going around a unit circle in the xy-plane a) counterclockwise from 0o to 180o b) clockwise from 0o to -180o, doing against a force field given by:

OpenStudy (anonymous):

\[F = \frac{ -i y }{ x^{2} + y^{2} } + \frac{ -j x }{ x^{2} + y^{2} }\] Note that the work done depends on the path

OpenStudy (anonymous):

@ivanmlerner @sandra

OpenStudy (anonymous):

\[W=-\int\limits_{x(t_{0})}^{x(t_{1})}F.rdl\]Use cilindrical coordinates and you get:\[F=\frac{-\sin(\theta)}{r^2}i+\frac{-\cos(\theta)}{r^2}j\]\[dl=r d \theta\]Since r=1,

OpenStudy (anonymous):

Sorry, didnt finish, clicked post without wanting to, just a sec

OpenStudy (anonymous):

And its r, not r^2 on the denominator

OpenStudy (anonymous):

\[F.r=\left| F \right|\cos \left( 2\theta+\frac{\pi}{2}\right)\]From now on its just math.

OpenStudy (anonymous):

Do you understand it?